A driver operating a vehicle at a speed of 25.0 m/
Answer #1

14) time in the first case = distance travelled by second vehicle / its speed

= 2km/(30m/s) = 66.67s

using second equation of motion for first vehicle,

s = ut + 0.5at^2

4000 = 25*66.67 + 0.5*a*66.67^2

a = 2333/[0.5*66.67^2]

= 1.05 m/s^2

now in second case, let they both meet at t

distance travelled by first vehical = distance travelled by first vehical + 2000

25t + 0.5*1.05 t^2 = 30t + 0.5*0.4*t^2 +2000

(0.5*1.05-0.5*0.4) t^2 - 5t -2000 = 0

0.325 t^2 -5t -2000 =0

solving the quadratic equation,

t =86 s, and a=1.05 should be correct answer of 14 and 15 respectively [checked many times]

However, among the given options, following is correct

Answer of Q14= B, 50s

Answer of Q15 = C 2.2m/s^2